OpenStudy (meilendurcer):

A recursive function is shown below: f(1) = 5 and f(n) = f(n - 1) - 6; n > 1 Which of the following lists the terms in the sequence defined by this recursive function? A. 5, 1, 7, 13, 19, ... B. 5, -1, -7, -13, -19, ... C. 5, 11, 17, 23, 29, ... D. 5, -11, -17, -23, -29, ... I think the answer is A.

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Connections?

OpenStudy (meilendurcer):

nope

OpenStudy (anonymous):

im on connections

OpenStudy (loser66):

@The_Chosen A is not correct

OpenStudy (anonymous):

what would u know u are a loser

OpenStudy (meilendurcer):

@Loser66 that's what i thought after looking at it a while.B or C?

OpenStudy (loser66):

let say, if n =2 , then f(2) = f(1) -6 = 5-6 = -1, not 1 when you choose A, the second term is 1, so that's wrong

OpenStudy (meilendurcer):

ok i get it a little

OpenStudy (loser66):

Of course. I am a Loser (big Loser, not small loser). However, I am a Loser at my level, not yours. hehehehe

OpenStudy (loser66):

@meilendurcer only 1 option has the second term is -1. You pick it

OpenStudy (meilendurcer):

Tha snwer surely isn't D right? can i eliminate that?

OpenStudy (meilendurcer):

@loser66 ah k so B thanks! could you possibly help with more?